# Transparent Image Layer Diffusion using Latent Transparency

### v3

@@ 3.2 Diffusion Model with Latent Transparency
Figure 3: Model Training. We visualize the training of the base model to generate transparent images, and the training of the multi-layer model to generate multiple layers together. When training the base diffusion model (a), all model weights are trainable, whereas for training the multi-layer model (b), only two LoRAs are trainable (the foreground LoRA and background LoRA).
3.3 Generating Multiple Layers
We further extend the base model to a multi-layer model using attention sharing and LoRAs [17], as shown in Fig. 3-(b). We denote the foreground noisy latent as $\bm{x}_{f}$ and background as $\bm{x}_{b}$, and train two LoRAs, a foreground LoRA parameterized by $\theta_{\text{f}}$ and a background LoRA by $\theta_{\text{b}}$, to denoise the latent images. If the two models independently denoise the two images, we have the two objectives with
$\mathbb{E}_{\bm{x}_{f},t,\bm{c}_{t},\epsilon_{f}\sim\mathcal{N}(0,1)}&\Big{[}\|\epsilon_{f}-\epsilon_{\theta,\theta_{\text{f}}}(\bm{x}_{f},t,\bm{c}_{t}))\|_{2}^{2}\Big{]}\\ \mathbb{E}_{\bm{x}_{b},t,\bm{c}_{t},\epsilon_{b}\sim\mathcal{N}(0,1)}&\Big{[}\|\epsilon_{b}-\epsilon_{\theta,\theta_{\text{b}}}(\bm{x}_{b},t,\bm{c}_{t}))\|_{2}^{2}\Big{]}\end{aligned}\right.$ (8)
@@ 3.3 Generating Multiple Layers
$\mathcal{L}_{\text{layer}}=\mathbb{E}_{\bm{x}_{f},\bm{x}_{b},t,\bm{c}_{t},\epsilon_{m}\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon_{m}-\epsilon_{\theta,\theta_{\text{f}},\theta_{\text{g}}}(\bm{x}_{f},\bm{x}_{b},t,\bm{c}_{t}))\|_{2}^{2}\Big{]}$ (9)
@@
We also introduce several alternative architectures in Fig. 4 for more complicated workflows. We can add zero-initialized channels to the UNet and use VAE (with or without latent transparency) to encode foreground, or background, or layer combinations into conditions, and train the model to generate foreground or background (e.g., Fig. 4-(b, d)), or directly generate blended images (e.g., Fig. 4-(a, c)).
Figure 4: Alternative Architectures for More Complicated Workflows. We also include several alternative models for more complicated workflows. These include generating a blended image from background or foreground, as well as generating background/foreground from other combined layers.
3.4 Dataset Preparation and Training Details
@@
We use a human-in-the-loop method to collect a dataset of transparent images and train our models. The dataset initially contains 20k high-quality transparent PNG images purchased or downloaded free from 5 online image stocks (all images include commercial use permission (examples in Fig. 5-(a)). We then train the SDXL VAE with latent transparency using randomly sampled images with equal probability (at batch size 8), and then train the SDXL diffusion model using the same data with adjusted latents. Next we repeat the following steps for a total of 25 rounds. At the beginning of each round, we generate 10k random samples using the last model in the previous round. and the random prompts from LAIONPOP [38]. We then manually pick 1000 samples to add back to the training dataset. The newly added samples are given a 2x higher probability of appearing in training batches in the next round. We then train the latent transparency encoder-decoder and diffusion models again. After 25 rounds, the size of the dataset increases to 45K. Afterwards, we generate 5M sample pairs without human interaction and use the LAION Aesthetic threshold [39] setting of 5.5 and clip score sorting to obtain 1M sample pairs. We automatically remove samples that do not contain any transparent pixels as well as those that do not contain any visible pixels. Finally, all images are captioned with LLaVA [29] (an open-source multi-modal GPT similar to GPT4v) to get detailed text prompts. The training of both the VAE and the diffusion model is finalized with another 15k iterations using the final 1M dataset.
Figure 5: Dataset Preparation. We demonstrate the preparation of the two dataset: the transparent image dataset (base dataset) and multi-layer dataset. The base dataset is collected by downloading online transparent images with the help of a human-in-the-loop training method. The multi-layer dataset is synthesized with our transparent diffusion model and several state-of-the-art models including ChatGPT, SDXL inpaint model, etc. The final scale of each dataset is around 1M.
@@
We further extend our {text, transparent image} dataset into a {text, foreground layer, background layer} dataset, so as to train the multi-layer models. As shown in Fig. 5-(b), we ask GPTs (we used ChatGPT for 100k requests and then moved to LLAMA2 for 900k requests) to generate structured prompts pairs for foreground like “a cute cat”, entire image like “cat in garden”, and background like “nothing in garden” (we ask GPT to add the word “nothing” to the background prompt). The foreground prompt is processed by our trained transparent image generator (Section 3.2) to obtain the transparent images. Then, we use Diffusers Stable Diffusion XL Inpaint model [10] to inpaint all pixels with alpha less than one to obtain intermediate images using the prompt for the entire images. Finally, we invert the alpha mask, erode $k=8$ pixels and inpaint again with the background prompt to get the background layer. We repeat this process 1M times to generate 1M layer pairs.
@@
We use the AdamW optimizer at learning rate 1e-5 for both VAE and diffusion model. The pretrained Stable Diffusion model is SDXL [32]. For the the LoRA [17] training, we always use rank 256 for all layers. We use the Diffusers’ standard for naming and extracting LoRA keys. In the human-in-the-loop data collection, each round contains 10k iterations at batch size 16. The training devices are 4x A100 80G NV-link, and the entire training takes one week (to reduce budget, the training is paused when human are collection data for the next round of optimization) and the real GPU time is about 350 A100 hours. Our approach is training friendly for personal-scale or lab-scale research as the 350 GPU hours can often be processed within 1K USD.
4 Experiments
@@
We detail qualitative and quantitative experiments with our system. We first present qualitative results with single images (Section 4.1), multiple layers (Section 4.2), as well as iterative generation (Section 4.3), and then show that our framework can also be combined with control modules for wider applications (Section 4.4). We then analysis the importance of each component with ablative study (Section 4.5), and then discuss the difference and connection between our approach and image matting (Section 4.6). Finally, we conduct perceptual user study (Section 4.7) and present a range of discussions to further study the behaviors of our framework (Section 4.8,4.9,4.10).
4.1 Qualitative Results
”glass cup” ”man, 4k, best-quality” ”animal” ”game assets with magic effects” Refer to caption
Figure 6: Qualitative Results. We showcase various examples of transparent images generated by our model. The prompts for each group is given at the top of the examples. These examples only use our base single-layer model.
Refer to caption ”dragon over castle” ”dog in garden” ”woman, messy hair” ”man in room” ”robot in street” ”apple on table”
Figure 7: Multi-Layer Qualitative Results. We presents qualitative results generated by our model using prompts with diverse topics. For each example, we show the blended image, and two output layers. More results are available in supplementary materials.
@@ 4.1 Qualitative Results
Figure 8: Conditional Layer Generating. We presents results with foreground-conditioned background (the first two rows) and background-conditioned foreground (the last two rows). For each example, we generate two foregrounds/backgrounds.
@@
Figure 9: Generating Multiple Layers. We show that our framework can compose multiple layers iteratively, by repeating the background-conditioned foreground model. At each step, we blend al existing layers and feed the blended result to the background-conditioned generator. The prompts at each step is at the bottom of outputs.
@@
Figure 10: Combining with Control Models. We show that our approach can directly be combined with control models like ControlNet [59] to enhance the functionality. The prompts are “human in street”, “human in forest”, “big reflective ball in street”, and “big reflective ball in forest”.
@@
Figure 11: Ablative Study. We compare our approach to two alternative architecture: directly adding channels to UNet and directly adding channels to VAE. When adding channel to UNet, we directly encode alpha channel as an external image and add 4 channels to UNet. When adding channels to VAE, the UNet is finetuned on the latent images encoded by the newer VAE. The test prompts are “fox”, “elder woman”, “a book”, “man”.
Refer to caption Blended (ours) Background (ours) Foreground (ours) [9] [26] User trim map [57]
@@
Figure 12: Difference between Joint Layer Generating and Generating-then-matting. This is not a result comparison since the left-most images are outputs, and our method does not decompose images. We presents our jointly generated layers on the left and try to reproduce similar results using matting approaches. The prompts are “fire on burning wood in forest”, “white fox in white snow ground, all white, very white”, and “basketball”.
@@
4.2 Conditional Layer Generation
@@ 4.3 Iterative Generation
4.4 Controllable Generation
As shown in Fig. 10, we demonstrate that existing control models like ControlNet [59] can be applied to our model for enriched functionality. We can see that the model is able to preserve the global structure according to the ControlNet signal to generate harmonious compositions with consistent illumination effects. We also use a “reflective ball” example to show that the model is able to interact with the content of the foreground and background to generate consistent illumination like the reflections.
4.5 Ablative Study
We conduct an ablative study to evaluate the contribution of each component in our framework. We are interested in a possible architecture that does not modify Stable Diffusion’s latent VAE encoder/decoder, but only add channels to the UNet. In the original Stable Diffusion, a $512\times 512\times 3$ image is encoded to a latent image of size $64\times 64\times 4$. This indicates that if we duplicate the $512\times 512\times 1$ alpha channel 3 times into a $512\times 512\times 3$ matrix, the alpha could be directly encoded into a $64\times 64\times 4$ latent image. By concatenating this with the original latent image, the final latent image would form a a $64\times 64\times 8$ matrix. This means we could add 4 channels to Stable Diffusion UNet to force it support an alpha channel. We present the results of this approach in Fig. 11-(a). We can see that this method severely degrades the generation quality of the pretrained large model, because its latent distribution is changed; although the VAE is unchanged (it is frozen), the additional 4 channels significantly change the feature distribution after the first convolution layer in the VAE UNet. Note that this is different from adding a control signal to the UNet — the UNet must generate and recognize the added channels all at the same time because diffusion is a iterative process, and the outputs of any diffusion step become the input of the next diffusion step.
@@ 4.5 Ablative Study
4.6 Relationship to Image Matting
We discuss the difference and connection between native transparent image generation and image matting. To be specific, we test the following matting methods: (1) PPMatting [9] is a state-of-the-art neural network image matting model. This model reports to achieve the highest precision among all “classic” neural network based matting methods, i.e., neural models trained from scratch on a collected dataset of transparent images. This model is fully automatic and does not need a user-specified tri-map. (2) Matting Anything [26] is a new type of image matting model based on the recently released Segment Anything Model (SAM) [22]. This model uses pretrained SAM as a base and finetunes it to perform matting. This model also does not need a user-specified tri-map. We also include a tri-map-based method to study the potential for user-guided matte extraction. (3) VitMatte [57] is a state-of-the-art matting model that uses trim maps. The architecture is a Vision Transformer (ViT) and represents the highest quality of current user-guided matting models.
Table 1: User Study. We present the results from user study. We conduct user study in two groups: the first group compares outputs between different methods, while the second group directly compare our generated results to the search result of a commercial transparent image assets (Adobe Stock). Higher is better and best in bold.
@@ 4.6 Relationship to Image Matting
SD + PPMatting [9] 2.1$\pm$1.2/
SD + Matting Anything [26] 0.8$\pm$0.5/
@@
4.7 Perceptual User Study
Figure 13: Raw outputs of the RGB channels and alpha channel. We present the raw RGB and alpha channel for evaluation. The prompts are “woman with messy hair”, “boy with messy hair”, and “glass cup”.
In order to perceptually evaluate and compare our approach with existing methods, we perform a perceptual user study focusing on human aspects of our native transparent results and ad-hoc method like Stable Diffusion + generation-and-matting. In particular, the user study involves 14 individuals, where 11 individuals are online crowd-source workers, 1 is a computer science student, and the other 2 are professional content creators. We sample 100 results using 3 methods (prompts are randomly sampled from PickaPic [23]), and this leads to 100 result groups, with each group containing 3 results from 3 methods. The participants are invited to rank the results in each group. When ranking the results in each group, we ask users the question – “Which of the following results do you prefer most? Please rank the following transparent elements according to your preference”. We use the preference rate as the testing metric. This process is repeated 4 times to compute the standard deviation. Afterwards, we calculate the average preference rate of each method. We call this user study “group 1”.
We compare our approach with SD+[9], SD+[26]. Here in, the “SD+” means first using Stable Diffusion XL to generate a RGB image, and then performing matting. Results are shown in Table. 1, group 1. We find that users prefer our approach over all other approaches (in more than 97% cases). This demonstrates the advantage of native transparent image generation over ad-hoc solution like generation-then-matting.
We also perform another simple user preference experiment, “group 2”, to compare our results against the searching for commercial transparent assets from Adobe Stock, using the same aforementioned user preference metric. In Table. 1, group 2, we report that the preference rate of our method is very close to commercial stock (45.3% v.s. 54.7%). Though the high-quality paid content from commercial stock is still preferred marginally. This result indicates that our generated transparent content is competitive to commercial sources that require users to pay for each image.
4.8 Raw RGBA Channels

### v4

@@ 3.2. Diffusion Model with Latent Transparency
3.3. Generating Multiple Layers
Figure 4. Dataset Preparation. We demonstrate the preparation of the two datasets: the transparent image dataset (base dataset) and multi-layer dataset. The base dataset is collected by downloading online transparent images and a human-in-the-loop training method. The multi-layer dataset is synthesized with our transparent diffusion model and several state-of-the-art models including ChatGPT, SDXL inpaint model, etc. The final scale of each dataset is around 1M.
Figure 5. Human-in-the-loop data screening. We visualize sample examples that are preserved versus removed in each round during the dataset collection process. We show examples from the round 1, 5, 10, and 20. The prompts are randomly sampled during the collecting process.
”glass cup” ”man, 4k, best-quality” ”animal” ”game assets with magic effects”
Figure 6. Qualitative Results. We showcase various examples of transparent images generated by our model. The prompts for each group is given at the top of the examples. These examples only use our base single-layer model.
”dragon over castle” ”dog in garden” ”woman, messy hair” ”man in room” ”robot in street” ”apple on table”
Figure 7. Multi-Layer Qualitative Results. We presents qualitative results generated by our model using prompts with diverse topics. For each example, we show the blended image, and two output layers. More results are available in supplementary materials.
We further extend the base model to a multi-layer model using attention sharing and LoRAs (Hu et al., 2021), as shown in Fig. 3-(b). We denote the foreground noisy latent as $\bm{x}_{f}$ and background as $\bm{x}_{b}$, and train two LoRAs, a foreground LoRA parameterized by $\theta_{\text{f}}$ and a background LoRA by $\theta_{\text{b}}$, to denoise the latent images. If the two models independently denoise the two images, we have the two objectives with
(8) $\mathbb{E}_{\bm{x}_{f},t,\bm{c}_{t},\epsilon_{f}\sim\mathcal{N}(0,1)}&\Big{[}\|\epsilon_{f}-\epsilon_{\theta,\theta_{\text{f}}}(\bm{x}_{f},t,\bm{c}_{t}))\|_{2}^{2}\Big{]}\\ \mathbb{E}_{\bm{x}_{b},t,\bm{c}_{t},\epsilon_{b}\sim\mathcal{N}(0,1)}&\Big{[}\|\epsilon_{b}-\epsilon_{\theta,\theta_{\text{b}}}(\bm{x}_{b},t,\bm{c}_{t}))\|_{2}^{2}\Big{]}\end{aligned}\right.$
@@ 3.3. Generating Multiple Layers
(9) $\mathcal{L}_{\text{layer}}=\mathbb{E}_{\bm{x}_{f},\bm{x}_{b},t,\bm{c}_{t},\epsilon_{m}\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon_{m}-\epsilon_{\theta,\theta_{\text{f}},\theta_{\text{g}}}(\bm{x}_{f},\bm{x}_{b},t,\bm{c}_{t}))\|_{2}^{2}\Big{]}$
@@
3.4. Dataset Preparation and Training Details
@@
We use a human-in-the-loop method to collect a dataset of transparent images and train our models. The dataset initially contains 20k high-quality transparent PNG images purchased or downloaded free from 5 online image stocks (all images include commercial use permission (examples in Fig. 4-(a)). We then train the SDXL VAE with latent transparency using randomly sampled images with equal probability (at batch size 8), and then train the SDXL diffusion model using the same data with adjusted latents. Next we repeat the following steps for a total of 25 rounds. At the beginning of each round, we generate 10k random samples using the last model in the previous round. and the random prompts from LAIONPOP (Schuhmann and Bevan, 2023). We then manually pick 1000 samples to add back to the training dataset. The newly added samples are given a 2x higher probability of appearing in training batches in the next round. We then train the latent transparency encoder-decoder and diffusion models again. After 25 rounds, the size of the dataset increases to 45K. Afterwards, we generate 5M sample pairs without human interaction and use the LAION Aesthetic threshold (Schuhmann et al., 2022) setting of 5.5 and clip score sorting to obtain 1M sample pairs. We automatically remove samples that do not contain any transparent pixels as well as those that do not contain any visible pixels. Finally, all images are captioned with LLaVA (Liu et al., 2023) (an open-source multi-modal GPT similar to GPT4v) to get detailed text prompts. The training of both the VAE and the diffusion model is finalized with another 15k iterations using the final 1M dataset.
Table 1. Statistical Record of Human-in-the-loop Collection. We report the Defective Sample Count (DSC) per 100 sampling during the rounds of human data selection. We resume the model checkpoints recorded after each round of data collection and generate 100 samples for each checkpoint. Users find how many samples are of obvious defects (like fully empty image, or fully non-transparent image, or obvious errors like opaque glass, etc) and report the number as *DSC* (lower is better $\downarrow$).
Round #0 #1 #2 #3 #4 #5 #6 #7 #8 #9 #10 #12 #14 #16 #18 #20
DSC $\downarrow$ 61 62 37 53 41 25 17 15 23 21 25 11 6 9 3 5
Statistical Analysis
We briefly analyze here how human data selection improves the quality of the dataset as well as the model capabilities. As shown in Fig. 5, we visualize samples that are preserved or removed in each round of the human-in-the-loop selection. We can see that human efforts removes some obvious flaws (e.g., empty images, fully opaque colors for glass, etc.) and enhances the diversity of the dataset content (e.g., the glowing effects on magic books, etc.). In Table 1, we resume the checkpoint from each round of data collection to sample images, and ask the users to review the images and count images with obvious defects. We can see that as the number of rounds increases, the rate of defective outputs gradually decreases.
@@
We further extend our {text, transparent image} dataset into a {text, foreground layer, background layer} dataset, so as to train the multi-layer models. As shown in Fig. 4-(b), we ask GPTs (we used ChatGPT for 100k requests and then moved to LLAMA2 for 900k requests) to generate structured prompts pairs for foreground like “a cute cat”, entire image like “cat in garden”, and background like “nothing in garden” (we ask GPT to add the word “nothing” to the background prompt). The foreground prompt is processed by our trained transparent image generator (Section 3.2) to obtain the transparent images. Then, we use Diffusers Stable Diffusion XL Inpaint model (diffusers, 2024) to inpaint all pixels with alpha less than one to obtain intermediate images using the prompt for the entire images. Finally, we invert the alpha mask, erode $k=8$ pixels and inpaint again with the background prompt to get the background layer. We repeat this process 1M times to generate 1M layer pairs.
@@
We use the AdamW optimizer at learning rate 1e-5 for both VAE and diffusion model. The pretrained Stable Diffusion model is SDXL (Podell et al., 2023). For the the LoRA (Hu et al., 2021) training, we always use rank 256 for all layers. We use the Diffusers’ standard for naming and extracting LoRA keys. In the human-in-the-loop data collection, each round contains 10k iterations at batch size 16. The training devices are 4x A100 80G NV-link, and the entire training takes one week (to reduce budget, the training is paused when human are collection data for the next round of optimization) and the real GPU time is about 350 A100 hours. Our approach is training friendly for personal-scale or lab-scale research as the 350 GPU hours can often be processed within 1K USD.
4. Experiments
@@
We detail qualitative and quantitative experiments with our system. We first present qualitative results with single images (Section 4.1), multiple layers (Section 4.2), as well as iterative generation (Section 4.3), and then show that our framework can also be combined with control modules for wider applications (Section 4.4). We then analysis the importance of each component with ablative study (Section 4.5), and then discuss the difference and connection between our approach and image matting (Section 4.6). Finally, we conduct perceptual user study (Section 4.7) and present a range of discussions to further study the behaviors of our framework (Section 4.8,4.10,4.12).
4.1. Qualitative Results
@@
Figure 8. Conditional Layer Generating. We presents results with foreground-conditioned background (the first two rows) and background-conditioned foreground (the last two rows). For each example, we generate two foregrounds/backgrounds.
@@
Figure 9. Generating Multiple Layers. We show that our framework can compose multiple layers iteratively, by repeating the background-conditioned foreground model. At each step, we blend al existing layers and feed the blended result to the background-conditioned generator. The prompts at each step is at the bottom of outputs.
@@
Figure 10. Combining with Control Models. We show that our approach can directly be combined with control models like ControlNet (Zhang and Agrawala, 2023) to enhance the functionality. The prompts are “human in street”, “human in forest”, “big reflective ball in street”, and “big reflective ball in forest”.
@@
Figure 11. Ablative Study. We compare our approach to two alternative architecture: directly adding channels to UNet and directly adding channels to VAE. When adding channel to UNet, we directly encode alpha channel as an external image and add 4 channels to UNet. When adding channels to VAE, the UNet is finetuned on the latent images encoded by the newer VAE. The test prompts are “fox”, “elder woman”, “a book”, “man”.
Figure 12. Additional Ablative Architectures We also include several alternative models for more complicated workflows. These include generating a blended image from background or foreground, as well as generating background/foreground from other combined layers. We also demonstrate the use of two-step pipelines to generate/decompose independent layers.
Foreground (ours) Background (ours) Blended (ours) (Chen et al., 2022) (Li et al., 2023b) User tri-map (Yao et al., 2024)
@@
Figure 13. Difference between Joint Layer Generating and Generating-then-matting. This is not a result comparison since the left images are outputs layers of our method. The blended images are alpha blending of the generated layers. (Our method does not decompose images.) We try to reproduce similar results using matting approaches. The prompts are “fire on burning wood in forest”, “white fox in white snow ground, all white, very white”, and “basketball”.
@@
4.2. Conditional Layer Generation
@@ 4.3. Iterative Generation
4.4. Controllable Generation
As shown in Fig. 10, we demonstrate that existing control models like ControlNet (Zhang and Agrawala, 2023) can be applied to our model for enriched functionality. We can see that the model is able to preserve the global structure according to the ControlNet signal to generate harmonious compositions with consistent illumination effects. We also use a “reflective ball” example to show that the model is able to interact with the content of the foreground and background to generate consistent illumination like the reflections.
4.5. Ablative Study
We conduct an ablative study to evaluate the contribution of each component in our framework. We are interested in a possible architecture that does not modify Stable Diffusion’s latent VAE encoder/decoder, but only adds channels to the UNet. In the original Stable Diffusion, a $512\times 512\times 3$ image is encoded to a latent image of size $64\times 64\times 4$. This indicates that if we duplicate the $512\times 512\times 1$ alpha channel 3 times into a $512\times 512\times 3$ matrix, the alpha could be directly encoded into a $64\times 64\times 4$ latent image. By concatenating this with the original latent image, the final latent image would form a a $64\times 64\times 8$ matrix. This means we could add 4 channels to Stable Diffusion UNet to force it support an alpha channel. We present the results of this approach in Fig. 11-(a). We can see that this method severely degrades the generation quality of the pretrained large model, because its latent distribution is changed; although the VAE is unchanged (it is frozen), the additional 4 channels significantly change the feature distribution after the first convolution layer in the VAE UNet. Note that this is different from adding a control signal to the UNet — the UNet must generate and recognize the added channels all at the same time because diffusion is a iterative process, and the outputs of any diffusion step become the input of the next diffusion step.
@@ 4.5. Ablative Study
We also introduce several alternative architectures in Fig. 12 for more complicated workflows. We can add zero-initialized channels to the UNet and use the VAE (with or without latent transparency) to encode the foreground, or background, or layer combinations into conditions, and train the model to generate foreground or background or directly generate blended images (e.g., Fig. 12-(a, b, c)). We visualize examples of this two-stage pipeline in Fig. 12-(d, e).
4.6. Relationship to Image Matting
We discuss the difference and connection between native transparent image generation and image matting. To be specific, we test the following matting methods: (1) PPMatting (Chen et al., 2022) is a state-of-the-art neural network image matting model. This model reports to achieve the highest precision among all “classic” neural network based matting methods, i.e., neural models trained from scratch on a collected dataset of transparent images. This model is fully automatic and does not need a user-specified tri-map. (2) Matting Anything (Li et al., 2023b) is a new type of image matting model based on the recently released Segment Anything Model (SAM) (Kirillov et al., 2023). This model uses pretrained SAM as a base and finetunes it to perform matting. This model also does not need a user-specified tri-map. We also include a tri-map-based method to study the potential for user-guided matte extraction. (3) VitMatte (Yao et al., 2024) is a state-of-the-art matting model that uses tri-maps. The architecture is a Vision Transformer (ViT) and represents the highest quality of current user-guided matting models.
Table 2. User Study. We present the results from user study. We conduct user study in two groups: the first group compares outputs between different methods, while the second group directly compare our generated results to the search result of a commercial transparent image assets (Adobe Stock). Higher is better and best in bold.
@@ 4.6. Relationship to Image Matting
SD + PPMatting (Chen et al., 2022) 2.1$\pm$1.2/
SD + Matting Anything (Li et al., 2023b) 0.8$\pm$0.5/
@@
4.7. Perceptual User Study
Figure 14. Raw outputs of the RGB channels and alpha channel. We present the raw RGB and alpha channel for evaluation. The prompts are “woman with messy hair”, “boy with messy hair”, and “glass cup”.
w/ latent offset w/o latent offset + data augmentation
Figure 15. Robust decoder with data augmentations. We show that it is possible to use data augmentation methods to train a robust decoder to handle situations when the UNet cannot diffuse the desired latent offsets. Samples are transparent images on black backgrounds.
In order to perceptually evaluate and compare our approach with existing methods, we perform a perceptual user study focusing on human aspects of our native transparent results and ad-hoc methods like Stable Diffusion + generation-and-matting. We target real-world use cases where users want to get transparent elements given specific demands (prompts). Our study tests multiple types of methods (native transparent generation, generating-and-matting, online commercial stock) to see how they fulfill such demands (by asking users which they prefer).
Specifically, our user study involves 14 individuals, where 11 individuals are online crowd-source workers, 1 is a computer science student, and the other 2 are professional content creators. We sample 100 results using the 3 methods (prompts are randomly sampled from PickaPic (Kirstain et al., 2023)), and this leads to 100 result groups, with each group containing 3 results from 3 methods. The participants are invited to rank the results in each group. When ranking the results in each group, we ask users the question – “Which of the following results do you prefer most? Please rank the following transparent elements according to your preference”. We use the preference rate as the testing metric. This process is repeated 4 times to compute the standard deviation. Afterwards, we calculate the average preference rate of each method. We call this user study “group 1”.
We compare our approach with SD+PPMatting (Chen et al., 2022), SD+Matting Anything (Li et al., 2023b). Herein, “SD+” means we first use Stable Diffusion XL to generate an RGB image, and then perform matting using the corresponding method. Results are shown in Table. 2, group 1. We find that users prefer our approach over all other approaches (in more than 97% cases). This demonstrates the advantage of native transparent image generation over ad-hoc solution like generation-then-matting.
We also perform another user preference experiment in “group 2”, comparing our results against searching for commercial transparent assets from Adobe Stock, using the same aforementioned user preference metric. In Table. 2, group 2, we report that the preference rate of our method is close to commercial stock (45.3% v.s. 54.7%). Though the high-quality paid content from commercial stock is still preferred marginally. This result suggests that our generated transparent content is competitive to commercial sources that require users to pay for each image.
4.8. Raw RGBA Channels